# UFO3: Weaving the Digital Agent Galaxy [Uncaptioned image]

### v1

@@ 10.5 RQ2: Orchestration and Adaptation
Added
Tasks
Removed
Tasks
Modified
Tasks
Added
Dependencies
Removed
Dependencies
Modified
Dependencies
Total
@@ 10.6 RQ3: Parallelism and Execution Efficiency
Difficulty
Max
Parallel Width
Critical Path
Length (L)
Total Exec.
Time (W)
Parallelism
Ratio (P = W/L)
@@ 11.1 Scalability and Ecosystem Growth
Moreover, UFO3 orchestrates at the protocol level rather than binding to specific implementations. Any agent compatible with AIP, whether it is a coding agent yang2024swe, a deep-research agent huang2025deep, or a data-analysis module zhang2025allhands, can participate in the orchestration. This design opens the door for a broader agent ecosystem, where specialized agents cooperate under a unified orchestration fabric. Our long-term vision is to make UFO3 an “agent galaxy”, a shared runtime that interconnects diverse autonomous agents into a cohesive, cross-platform execution universe.
@@ 11.2 Limitations of Device-Centric Execution
At the core of UFO3 lies the ConstellationAgent, which governs dynamic DAG scheduling and reasoning. However, the system’s overall capability remains constrained by the competence of individual device agents. Since execution is distributed, a single agent’s failure, such as incomplete subtask execution or incorrect result reporting, can cascade through the constellation and lead to global task failure cemri2025multi. This “weakest-link effect” underscores the need for more reliable and self-verifying device agents. Future work will explore automatic capability validation, task-level sandboxing, and adaptive reallocation strategies that allow the orchestrator to detect and recover from unreliable nodes without compromising the consistency of the global task state.
@@ 11.3 Toward Shared Cross-Device Memory
Currently, agent communication in UFO3 is purely textual, suitable for command exchange but insufficient for data-rich workflows involving images, logs, or binaries. Many cross-device tasks, however, require file transfer or shared context, for example, a Windows agent generating a dataset for a Linux node to process. To support such workflows, we plan to extend AIP with a persistent, shared “agent memory” zhang2025survey, effectively a distributed workspace where agents can store and retrieve intermediate results. This enhancement would unify context across devices, simplify file sharing, and enable richer collaborative behavior across heterogeneous agents.
@@ 12.1 Digital Device Agents
The emergence of large language models, particularly multimodal variants, has enabled a new class of intelligent agents capable of operating directly on digital devices zhang2024large. These agents extend beyond web and mobile environments to desktops, and are now expanding into domains such as automotive and embedded systems.
Early systems like SeeAct zheng2024gpt, Mobile-Agent wang2024mobile, and UFO zhang2025ufo pioneered this direction by leveraging GPT-4V yang2023dawn to perceive graphical user interfaces and perform human-like interactions across different platforms. Subsequent research has evolved along two main trajectories, namely (i) system-level integration, enhancing robustness through tighter coupling with native OS and API layers zhang2025api; zhang2025ufo2; and (ii) model-level adaptation, fine-tuning LLMs with large-scale interaction data and reinforcement learning to improve grounding and autonomy in digital environments luo2025gui; zheng2025vem; wang2025ui.
Industry adoption has accelerated rapidly. Anthropic’s Claude (Computer Use) anthropic2024 relies purely on screenshot-based perception, while OpenAI’s Operator cua2025 demonstrates strong multimodal reasoning and robust desktop automation. These advances collectively signal the formation of a new computing paradigm where LLMs serve as universal control interfaces for heterogeneous software systems.
@@ 12.2 Multi-Agent Orchestration
Recent years have witnessed growing interest in LLM-based multi-agent systems, where coordination among heterogeneous agents becomes central to achieving emergent intelligence guo2024large. Designing robust orchestration protocols, covering role assignment, communication, collaboration, and context sharing, has been identified as a key challenge bhatt2025should; dang2025multi; kong2025survey.
The Internet of Agents (IoA) cheninternet represents one of the earliest structured attempts in this direction. It introduces an integration protocol and an instant-messaging-style communication layer, enabling flexible, scalable multi-agent collaboration. IoA emphasizes dynamic teaming and conversation flow control, allowing agents to self-organize and cooperate fluidly in an Internet-like environment yang2025agentic. Building upon this idea, the Federation of Agents (FoA) giusti2025federation further advances distributed orchestration with Versioned Capability Vectors (VCVs) as machine-readable profiles that encode agent capabilities, costs, and constraints. These semantic embeddings make capabilities discoverable and composable, enabling dynamic, capability-driven coordination at scale.

### v2

@@ 10.5 RQ2: Orchestration and Adaptation
Added Tasks Removed Tasks Modified Tasks Added Dependencies Removed Dependencies Modified Dependencies Total
@@ 10.6 RQ3: Parallelism and Execution Efficiency
Difficulty Max Parallel Width Critical Path Length (L) Total Exec. Time (W) Parallelism Ratio (P = W/L)
@@ 11.1 Scalability and Ecosystem Growth
Moreover, UFO3 orchestrates at the protocol level rather than binding to specific implementations. Any agent compatible with AIP, whether it is a coding agent Yang et al. (2024), a deep-research agent Huang et al. (2025), or a data-analysis module Zhang et al. (2025c), can participate in the orchestration. This design opens the door for a broader agent ecosystem, where specialized agents cooperate under a unified orchestration fabric. Our long-term vision is to make UFO3 an “agent galaxy”, a shared runtime that interconnects diverse autonomous agents into a cohesive, cross-platform execution universe.
@@ 11.2 Limitations of Device-Centric Execution
At the core of UFO3 lies the ConstellationAgent, which governs dynamic DAG scheduling and reasoning. However, the system’s overall capability remains constrained by the competence of individual device agents. Since execution is distributed, a single agent’s failure, such as incomplete subtask execution or incorrect result reporting, can cascade through the constellation and lead to global task failure Cemri et al. (2025). This “weakest-link effect” underscores the need for more reliable and self-verifying device agents. Future work will explore automatic capability validation, task-level sandboxing, and adaptive reallocation strategies that allow the orchestrator to detect and recover from unreliable nodes without compromising the consistency of the global task state.
@@ 11.3 Toward Shared Cross-Device Memory
Currently, agent communication in UFO3 is purely textual, suitable for command exchange but insufficient for data-rich workflows involving images, logs, or binaries. Many cross-device tasks, however, require file transfer or shared context, for example, a Windows agent generating a dataset for a Linux node to process. To support such workflows, we plan to extend AIP with a persistent, shared “agent memory” Zhang et al. (2025f), effectively a distributed workspace where agents can store and retrieve intermediate results. This enhancement would unify context across devices, simplify file sharing, and enable richer collaborative behavior across heterogeneous agents.
@@ 12.1 Digital Device Agents
The emergence of large language models, particularly multimodal variants, has enabled a new class of intelligent agents capable of operating directly on digital devices Zhang et al. (2024). These agents extend beyond web and mobile environments to desktops, and are now expanding into domains such as automotive and embedded systems.
Early systems like SeeAct Zheng et al. (2024), Mobile-Agent Wang et al. (2024a), and UFO Zhang et al. (2025b) pioneered this direction by leveraging GPT-4V Yang et al. (2023) to perceive graphical user interfaces and perform human-like interactions across different platforms. Subsequent research has evolved along two main trajectories, namely (i) system-level integration, enhancing robustness through tighter coupling with native OS and API layers Zhang et al. ; Zhang et al. (2025a); and (ii) model-level adaptation, fine-tuning LLMs with large-scale interaction data and reinforcement learning to improve grounding and autonomy in digital environments Luo et al. (2025); Zheng et al. (2025); Wang et al. (2025).
Industry adoption has accelerated rapidly. Anthropic’s Claude (Computer Use) Anthropic (2024) relies purely on screenshot-based perception, while OpenAI’s Operator OpenAI (2025a) demonstrates strong multimodal reasoning and robust desktop automation. These advances collectively signal the formation of a new computing paradigm where LLMs serve as universal control interfaces for heterogeneous software systems.
@@ 12.2 Multi-Agent Orchestration
Recent years have witnessed growing interest in LLM-based multi-agent systems, where coordination among heterogeneous agents becomes central to achieving emergent intelligence Guo et al. (2024). Designing robust orchestration protocols, covering role assignment, communication, collaboration, and context sharing, has been identified as a key challenge Bhatt et al. (2025); Dang et al. (2025); Kong et al. (2025).
The Internet of Agents (IoA) Chen et al. represents one of the earliest structured attempts in this direction. It introduces an integration protocol and an instant-messaging-style communication layer, enabling flexible, scalable multi-agent collaboration. IoA emphasizes dynamic teaming and conversation flow control, allowing agents to self-organize and cooperate fluidly in an Internet-like environment Yang et al. (2025b). Building upon this idea, the Federation of Agents (FoA) Giusti et al. (2025) further advances distributed orchestration with Versioned Capability Vectors (VCVs) as machine-readable profiles that encode agent capabilities, costs, and constraints. These semantic embeddings make capabilities discoverable and composable, enabling dynamic, capability-driven coordination at scale.